OpenStudy (anonymous):

Find f[g(x)] if f(x) = x^2 + 1 and g(x) = x^5. a) (x^5 + 1)2 b) (x^2 + 1)5 c) x^10 + 1 d) None of these

OpenStudy (anonymous):

@Michele_Laino @Compassionate @ganeshie8 @eliassaab

OpenStudy (anonymous):

@pooja195

OpenStudy (michele_laino):

please note that you have to evaluate this: \[f\left( {g\left( x \right)} \right) = {\left( {g\left( x \right)} \right)^2} + 1\] namely you have to set x = g(x) into the formula of f(x)

OpenStudy (anonymous):

so would it be (x^2+1)^5?

OpenStudy (anonymous):

then plug that back into x^2+1?

OpenStudy (michele_laino):

no, you have to solve this: \[f\left( {g\left( x \right)} \right) = {\left( {g\left( x \right)} \right)^2} + 1 = {\left( {{x^5}} \right)^2} + 1\]

OpenStudy (anonymous):

ohh okay x^10+1?:D

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

THANK U! :) once again amazing teacher!!!

OpenStudy (michele_laino):

Thank you! :)